OpenStudy (anonymous):

Solve the equation (x+3)/3 - (3x-12)/5 =1 The answer is 9 but how?

hero (hero):

(9+3)/3 - (3(9)-12)/5 =1 12/3 - (27-12)/5 = 1 4 - (15)/5 = 1 4 - 3 =1 1 = 1

OpenStudy (saifoo.khan):

u can insert the values to check the answer.